OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (2,5)

OpenStudy (anonymous):

Yeah there different questions

OpenStudy (anonymous):

y=3x+b plug in the point (2,5): 5=3(2)+b 5=6+b b=-1 therefore equation of the line is: y= 3x-1